please help me solve this problem: a square poster has two sides measuring 2 ft less than the sides of a square sign. if the difference between their areas is 36sq feet, find the lengths of the sides of the poster and the sign.
Answer by mananth(16946)

let side of square sign be x
area = x^2
..
square poster side = (x-2)
area = (x-2)^2
..
x^2-(x-2)^2= 36
x^2-(x^2-4x+4)=36
x^2-x^2+4x-4=36
4x-4=36
+4
4x-4+4 = 36+4
4x=40
/4
x= 10 feet. side of sign
10-2=8 feet = side of poster
...
m.ananth@hotmail.ca